# Enhanced entanglement scaling and area-law charge fluctuations in a non-Fermi liquid of composite fermions

### v1

@@ Acknowledgements.
$\kappa_{\mathrm{w}}^{(\alpha)}=\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}\frac{1}{2}\int_{\partial\Omega}\mathrm{d}S_{x}\int_{\partial\Gamma}\mathrm{d}S_{k}\,\frac{|\hat{\mathbf{n}}_{x}\cdot\hat{\mathbf{n}}_{k}|}{2\pi},$ (S7)
with a discrete sum over these modes. The flux factor $|\hat{n}_{x}\cdot\hat{n}_{k}|/2\pi$ needs to be replaced by $1/2\pi$, accounting for the rotational symmetry of the cap around the axis, and the double counting factor of $1/2$ becomes unnecessary. Then, the Renyi entropy for a spherical cap of angle $\theta$ is:
$S_{\alpha}=\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}\ln(R\sin\theta)\sum_{i=1}^{2n-1}\int_{\partial\Omega}\frac{\mathrm{d}S_{x}}{2\pi}=\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}(2\sqrt{N}-1)\sin\theta\ln(R\sin\theta),$ (S8)
@@
$S_{2}=\left(\frac{1}{4}-\frac{1}{8\sqrt{N}}\right)\lambda_{s}\ln\lambda_{s}\approx\frac{1}{4}\,\lambda_{s}\ln\lambda_{s},$ (S9)
@@
in agreement with the torus in the large-$N$ limit. Note that our derivation here does not capture finite-size effects related to the curvature of the sphere, which diminish in the limit $N\to\infty$.
SV Charge fluctuations
In the main text, we presented the results for charge fluctuations in a subsystem, $\Delta^{2}N_{A}$, normalized by $\lambda$ and plotted as a function of $\ln\lambda$. This was done to show that charge fluctuations are clearly suppressed when plotted in the same way as the entanglement entropy, hence there is no logarithmic enhancement like in ordinary Fermi liquids, where charge fluctuations mirror the behavior of entanglement entropy. In Fig. S4 we show the same data for $\Delta^{2}N_{A}$ but plotted on a linear $\lambda$ scale, which shows the area law scaling of $\Delta^{2}N_{A}$. From the linear fit to this data, we have extracted the intercept $\lambda\to 0$ in the insets of Fig. S4. This gives us the super-universal corner contribution, $4\pi^{2}b(\pi/2)$, expected to approach $\nu$ [33, 32, 31].
Figure S4: Charge fluctuations across the bipartition for CFL states at different filling factors. In contrast to free fermions (solid black line), the charge fluctuations in CFL states are strongly suppressed and obey area-law scaling (fits to area-law are plotted as colored dashed lines). LLL projection alters the area law coefficient but preserves the universal corner contribution. Table 1 shows the extracted corner contributions, which are proportional to the filling.
SVI SWAP algorithm and entropy decomposition
Following Refs. [30, 58], in our Monte Carlo evaluations of $S_{2}$ Rényi entropy we replace the partial trace with an expectation value $\langle\mathrm{SWAP}_{A}\rangle$ of the SWAP operator between two copies of the system, as illustrated in Fig. S5. This expectation value can be naturally broken down into a product of three terms:
$\langle\mathrm{SWAP}_{A}\rangle=P_{\mathrm{SWAP}}\,\langle\mathrm{SWAP}_{\mathrm{mod}}\rangle\langle\mathrm{SWAP}_{\mathrm{sign}}\rangle\,.$ (S10)
$P_{\mathrm{SWAP}}$ $=\frac{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,|\Psi(\alpha_{1},\beta_{1})|^{2}|\Psi(\alpha_{2},\beta_{2})|^{2}\delta_{N_{1},N_{2}}}{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,|\Psi(\alpha_{1},\beta_{1})|^{2}|\Psi(\alpha_{2},\beta_{2})|^{2}}\,,$ (S11)
Figure S5: Illustration of the SWAP process. We have two copies of the full system, whose particles are labeled with different colors, and each copy is divided into two regions separated by the dashed circle. Through the SWAP process, we interchange the particle coordinates inside the circle between the two copies while keeping the outside coordinates unchanged.
The following two terms in Eq. (S10) are the “mod” and “sign” contributions, which are evaluated only in the subspace of swappable copies ($N_{1}=N_{2}$),
$\langle\mathrm{SWAP}_{\mathrm{mod}}\rangle$ $=\frac{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{1},\beta_{1})|^{2}|\psi(\alpha_{2},\beta_{2})|^{2}\left|\frac{\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{2})}{\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})}\right|}{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{1},\beta_{1})|^{2}|\psi(\alpha_{2},\beta_{2})|^{2}},$ (S12)
$\langle\mathrm{SWAP}_{\mathrm{sign}}\rangle$ $=\frac{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{2})\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})|\,e^{i\theta(\bm{\alpha},\bm{\beta})}}{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{2})\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})|}\;.$ (S13)
$\exp i\theta(\bm{\alpha},\bm{\beta})=\frac{\psi(\alpha_{1},\beta_{2})^{*}\psi(\alpha_{2},\beta_{1})^{*}\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})}{|\psi(\alpha_{1},\beta_{2})\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})|}\;.$ (S14)
$S_{2}=S_{2,\mathrm{p}}+S_{2,\mathrm{mod}}+S_{2,\mathrm{sign}}.$ (S15)
In previous works, such as Ref. [28], the first two contributions to the entropy were lumped together, because $S_{2,\mathrm{p}}+S_{2,\mathrm{mod}}\equiv S_{2}(|\Psi|)$ can be interpreted as the entropy of the absolute value of the wave function $\Psi$. While this may be useful for gapped states [58], our data below suggests this is not helpful for gapless CFL states since $S_{2,\mathrm{p}}$ and $S_{2,\mathrm{mod}}$ exhibit different scaling with $\lambda$. Hence, we will study individually all three distinct contributions to $S_{2}$.
SVI.1 Mod and sign entropy for free fermions
The scaling of different contributions to entropy in Eq. (S15) is not rigorously known, even for free fermions where the full entropy has an established analytic expression. Fig. S6 contrasts the full $S_{2}$ entropy for free fermions [Fig. S6(a)] against its decompositions in Eq. (S15) obtained via Monte Carlo [Fig. S6(b)-(d)]. While the full entropy can be computed for large free-fermion systems containing thousands of particles, the evaluation of mod-sign entropy contributions, to the best of our knowledge, can only be done with Monte Carlo. This limits the accessible system sizes to $N\lesssim 100$, which is similar to the range of system sizes for the CFL states.
Within the range $N\lesssim 100$, we already observe the full entropy $S_{2}$ steadily converging to the Widom slope in Fig. S6(a). Figs. S6(b)-(c) illustrate the importance of considering the $\mathrm{p}$ and $\mathrm{mod}$ contributions to the entropy separately. They each converge nicely to the respective scalings, $S_{2,p}\sim\mathcal{O}(\ln\lambda)$ and $S_{2,\mathrm{mod}}\sim\mathcal{O}(\lambda)$. However, if we add them together to make up $S_{2}(|\Psi|)$, Fig. S6(c), the latter converges non-monotonically when plotted as $S_{2}(|\Psi|)/\lambda$ vs $\ln\lambda$. If we estimate the slope of $S_{2}(|\Psi|)$ by adding the fits, we obtain the solid line in Fig. S6(c), which suggests that our system sizes are not large enough to see the asymptotic area law scaling $S_{2}(|\Psi|)\sim\lambda$ that we expect in the thermodynamic limit. This has an impact on the scaling of the sign entropy in Fig. S6(d), which overshoots the Widom slope to compensate for the downturn in $S_{2}(|\Psi|)$. By plotting the Widom slope minus the extrapolation of $S_{2}(|\Psi|)$ using a solid line in Fig. S6(d), we see that it may only approach the Widom scaling in much larger system sizes ($\ln\lambda\gtrsim 4$). At the same time, the full entropy in Fig. S6(a) tracks the Widom scaling much more closely even in smaller system sizes. As the range of system sizes in this example is comparable to the CFL case, there appears to be little value in considering the mod and sign decompositions over the full entropy.
Figure S6: SWAP decomposition of the $S_{2}$ Rényi entropy for free fermions on the sphere. (a) Total $S_{2}$ entropy, computed via Monte Carlo (markers), while exact diagonalization result (solid line) is for the largest system size $N=625$. All system sizes steadily approach the Widom slope $\kappa_{\mathrm{w}}{=}0.25$ (dashed line). (b): $S_{2,\mathrm{p}}$ and $S_{2,\mathrm{mod}}$ that make up the absolute value entropy $S_{2}(|\Psi|)$. $S_{2,\mathrm{p}}$, given by the logarithm of the probability that the bipartitions of two copies of the system are swappable, scales logarithmically with the subsystem size. $S_{2,\mathrm{mod}}$ appears to follow an area-law growth. Black lines are extrapolations $\lambda\to\infty$. (c) The absolute value entropy $S_{2}(|\Psi|)=S_{2,p}+S_{2,\mathrm{mod}}$. The black line is the sum of the extrapolations in (b), which eventually settles to an area law for very large systems. However, unlike the individual decompositions in (b), their sum here appears far from the asymptotic scaling regime. (d) The sign entropy $S_{2,\mathrm{sign}}$. The dashed line is the Widom slope, while the continuous black line is the Widom slope minus the extrapolation of $S_{2}(|\Psi|)$, showing the difficulty in extracting the Widom slope solely from the sign term.
SVI.2 Mod and sign entropy for CFL
For completeness, in Fig. S7 we compare the full $S_{2}$ entropy against its mod-sign decompositions for the four CFL states considered ($\nu{=}1/m$, $m=1,2,3,4$). As mentioned in the main text, all of these cases display a significant enhancement of full entropy scaling. Mod-sign entropy decompositions, on the other hand, are plagued by similar issues as the free-fermion case discussed above. For example, in all the cases shown in Fig. S7, the mod contribution to the entropy is still quite large and, in fact, comparable to the sign entropy, which is expected to be the dominant contribution. Similarly, the expected area-law of $S_{2,\mathrm{mod}}$, i.e., $S_{2,\mathrm{mod}}/\lambda$ becoming flat when plotted as a function of $\ln\lambda$, is not yet reached in Fig. S7. This precludes drawing a clear conclusion based on the isolated sign entropy scaling. One new issue that arises in the CFL case is the large discrepancy between projected and unprojected CFL wave functions. This difference appears to be an artifact of the decomposition, since the full $S_{2}$ entropies are essentially unaffected by LLL projection, as we can see from the first panel in each row of Fig. S7.
Figure S7: Entropy decompositions at $\nu=1$ (first row), $\nu=1/2$ (second row), $\nu=1/3$ (third row), and $\nu=1/4$ (fourth row). Each row contains the data for both the projected and unprojected CFL wave functions on the torus. The first panel in each row shows the total entropy $S_{2}$, which is only affected by the projection in the case of $\nu=1$. The remaining three panels in each row show the p, mod, and sign contribution to the entropy.
SVI.3 Flux-attached wave functions
@@
$\Psi_{m}^{\mathrm{CFL,flux}}=\mathrm{Det}\big{[}\chi_{n}(z_{j})\big{]}{\prod}_{i{<}j}\left(\frac{(z_{i}{-}z_{j})}{|z_{i}-z_{j}|}\right)^{m}e^{-\frac{1}{4}\sum_{k}|z_{k}|^{2}}.$ (S16)
@@
Figure S8 shows results for $S_{2,\mathrm{sign}}$ for the state in Eq. (S16), with $m{=}2$ in the spherical geometry. We see that the sign entropy is significantly enhanced, even surpassing the standard vortex-attached wave functions discussed in the main text, which is in good agreement with the lattice realization of the flux-attached wave function [77]. Unfortunately, we are unable to efficiently project this wave function, as a significant proportion of it resides outside the lowest Landau level.
Figure S8: Difference in sign entropy $S_{2,\mathrm{sign}}-S_{2,\mathrm{sign}}^{\mathrm{free}}$ between the flux-attached CFL at filling $\nu{=}1/2$ and free fermions, on the sphere. We notice that the entanglement scaling of the sign entropy is considerably enhanced.

### v2

@@ Appendix E The Widom formula on the sphere
$\kappa_{\mathrm{w}}^{(\alpha)}=\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}\frac{1}{2}\int_{\partial\Omega}\mathrm{d}S_{x}\int_{\partial\Gamma}\mathrm{d}S_{k}\,\frac{|\hat{\mathbf{n}}_{x}\cdot\hat{\mathbf{n}}_{k}|}{2\pi},$ (23)
with a discrete sum over these modes. The flux factor $|\hat{n}_{x}\cdot\hat{n}_{k}|/2\pi$ needs to be replaced by $1/2\pi$, accounting for the rotational symmetry of the cap around the axis, and the double counting factor of $1/2$ becomes unnecessary. Then, the Rényi entropy for a spherical cap of angle $\theta$ is:
$S_{\alpha}$ $=$ $\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}\ln(R\sin\theta)\sum_{i=1}^{2n-1}\int_{\partial\Omega}\frac{\mathrm{d}S_{x}}{2\pi}$ (24)
$=$ $\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}(2\sqrt{N}-1)\sin\theta\ln(R\sin\theta),$
@@
$S_{2}=\left(\frac{1}{4}-\frac{1}{8\sqrt{N}}\right)\lambda_{s}\ln\lambda_{s}\approx\frac{1}{4}\,\lambda_{s}\ln\lambda_{s},$ (25)
@@
in agreement with the torus in the large-$N$ limit. Note that our derivation here does not capture finite-size effects related to the curvature of the sphere, which diminishes in the limit $N\to\infty$.
Appendix F Flux-attached wave functions
@@
$\Psi_{m}^{\mathrm{CFL,flux}}=\mathrm{Det}\big{[}\chi_{n}(z_{j})\big{]}{\prod}_{i{<}j}\left(\frac{(z_{i}{-}z_{j})}{|z_{i}-z_{j}|}\right)^{m}e^{-\frac{1}{4}\sum_{k}|z_{k}|^{2}}.$ (26)
@@
Figure 9: Difference in sign entropy $S_{2,\mathrm{sign}}-S_{2,\mathrm{sign}}^{\mathrm{free}}$ between the flux-attached CFL at filling $\nu{=}1/2$ and free fermions, on the sphere. We notice that the entanglement scaling of the sign entropy is considerably enhanced.
Figure 9 shows the results for $S_{2,\mathrm{sign}}$ of the state in Eq. (26), with $m=2$ in the spherical geometry. We see that the sign entropy is significantly enhanced, even surpassing the standard vortex-attached wave functions discussed in the main text, which is in good agreement with the lattice realization of the flux-attached wave function [82]. Unfortunately, we are unable to efficiently project this wave function as a significant proportion of it resides outside the lowest Landau level.